Two rainstorms occurred in one week in a certain area. The first storm lasted 25 hours, and the second storm lasted 40 hours, for a total 1600mL of rain. What was the rate of rainfall for each of the two storms if the sum of the two rates was 55 mL per hour?
Found 2 solutions by ptaylor, harecurtis:
Answer by ptaylor(2198)

Let x=rate of first rainstorm in ml per hour
Total rain from first rainstorm =25*x
Then 55-x=rate of the second rainstorm in ml per hour
Total rain from second rainstorm =40(55-x)
Now we know that the total rainfall (1600 ml) equals total rainfall from first storm (25x) plus total rainfall from second storm (40(55-x)), so our equation to solve is:
25x + 40(55-x)=1600 get rid of parens
25x+2200-40x=1600 subtract 2200 from each side
25x+2200-2200-40x=1600-2200 collect like terms
-15x=-600 divide both sides by -15
x=40 ml/hr----------------------------------rate of first rainstorm
55-x=55-40=15 ml/hr-------------------------rate of second rainstorm
CK
40*25+15*40=1600
1000+600=1600
1600=1600

Hope this helps--ptaylor



Answer by harecurtis(6)

Let x=rate of first rainstorm in ml per hour
Total rain from first rainstorm =25*x
Then 55-x=rate of the second rainstorm in ml per hour
Total rain from second rainstorm =40(55-x)
Now we know that the total rainfall (1600 ml) equals total rainfall from first storm (25x) plus total rainfall from second storm (40(55-x)), so our equation to solve is:
25x + 40(55-x)=1600 get rid of parens
25x+2200-40x=1600 subtract 2200 from each side
25x+2200-2200-40x=1600-2200 collect like terms
-15x=-600 divide both sides by -15
x=40 ml/hr----------------------------------rate of first rainstorm
55-x=55-40=15 ml/hr-------------------------rate of second rainstorm
CK
40*25+15*40=1600
1000+600=1600
1600=1600